# Constraint-Aware Feature Learning for Parametric Point Cloud

### v5

@@ 3.4 Constraint-Aware Feature Learning
The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD, Adj, PT separately, and attention between features is then applied to update constraint feature, with details in Fig. 7.
@@ 4.1 Experimental Settings
Param20K dataset Param20K contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes B-Rep data, mesh, and point cloud. All B-Rep data are collected from the following three sources. More details in the supplementary. Sources 1 (36.05 %): Downloaded and purified from TraceParts [18]. The CAD shapes on TraceParts are collected from manufacturing companies. Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components. Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
@@ 4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig. 8. By referring to the model row, CstBRep performs well in extracting MAD, Adj, and PT from B-Rep data. Predicting MAD, Adj, and PT from point cloud is achieved by the CstPnt module. The training set is sourced from 25 trunks of ABC [20], with each trunk containing 10,000 B-Rep files. We utilized CstBRep module to extract constraints as Ground Truth for model training. After the training finished, the MSE loss of MAD is 0.0559, the prediction accuracy of Adj and PT reached 94.62% and 94.23%, respectively. In subsequent experiments, the above pre-trained CstPnt module is utilized to predict MAD, Adj, and PT with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods when comparing, even when B-Rep data is available, the CstBRep module will not be activated to obtain MAD, Adj, and PT.
After CstPnt was pre-trained, we evaluated it on a ABC new trunk and Param20K dataset. The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available CAD shapes datasets in supplementary. Since there is no existing model that directly predicts MAD, Adj, and PT, we modified the published CAD-specific methods [58, 41] by adding MLPs as prediction head to enable these predictions. The experiment results are shown in Tab. 1, from which our CstPnt achieves the highest accuracy in predicting MAD, Adj, and PT. Furthermore, from ABC to Param20K, all methods’ performance decreased, but our method dropped least accuracy. ParSeNet and HPNet may be negatively affected by global features, which impaired their performance on unseen datasets. Fig. 8 presents the visualization of constraint prediction results, which shown that our method is closest to the ground truth. Furthermore, from ABC to Paeam20K, our method demonstrates excellent generalization performance on unseen datasets. In contrast, while ParSeNet and HPNet exhibit acceptable prediction results on the ABC new trunk, they performed unsatisfactory on the unseen Param20K dataset.
@@
Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in supplementary.
@@ 4.3 Classification
To validate the ability of our CstNet in identifying CAD shapes that are visual similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in the Problem Statement section Sec. 3.1. Results shown in Fig. 9, in which our CstNet achieved best instance accuracy in distinguish 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrate the strong ability of CstNet in recognizing shapes with similar appearance but constraints different. For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, this is attributed to the differences between them are so subtle that no one is capable to distinguishing them.
@@
The classification of prisms and cuboids is predominantly influenced by constraints, while the comprehension of common CAD shapes necessitates the integration of both constraint and geometric features. CstNet achieves this by concatenating XYZ with constraint in Stage 2, thereby integrating geometric features into constraint features. To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset (more results on public CAD shapes dataset in Supplementary), results are presented in Tab. 2, where our CstNet outperformed other methods across both metrics, validated the effectiveness of utilizing constraint and geometric features.
@@
To access the robustness of CstNet, we evaluated it under rotated point clouds. Since point clouds are pre-processed by aligning their centroids to the origin and scaling their ranges into a unit cube, the performance of all methods is unaffected by point cloud shifts and scaling. Comparison results depicted in Fig. 10, from which our CstNet demonstrates strongest robustness to point cloud rotation. The dashed line in Fig. 10 show that the performance of CstPnt module remains nearly unaffected by rotation, owing to it learned various local features from ABC dataset.
@@ 4.4 Ablation Studies
Our CstNet leverages all three components of constraint: MAD, Adj, and PT. Additionally, we utilized the SurfaceKNN to obtain more efficient neighbors for each point. To validate the effectiveness of our design, we conducted a series of ablation studies on Param20K dataset.
For each component of constraint, we evaluated their importance by disabling one or all at a time. This operation is performed by setting the disabled components as a zero tensor before feeding into CstNet. Classification results are presented in Tab. 3, which show that all three components contribute to improving the performance of our CstNet. To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 3, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
@@
To validate the effectiveness of Stage 2 backbone, we conducted a comprehensive evaluation by feeding (x, y, z, MAD, Adj, PT) into multiple classification backbone for comparison. The results are presented in Tab. 4, from which our method demonstrates best performance under both predicted and label constraints, indicating that our design achieves the best utilization of constraints. Furthermore, although the accuracy using predicted constraints is lower than that of using label constraints, it still shows a significant improvement compared to the baselines without constraints. This highlights the effectiveness of the predicted constraints in improving model performance.
Table 4: Effects on Stage2 Backbone. Upper half: predicted constraints from CstPnt, lower: label constraints.
@@ 5 Conclusions
In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly expression. Afterward, we proposed the CstNet for extracting and leveraging constraints. Additionally, we built up a multi-modal CAD shape dataset Param20K. Extensive experiment results demonstrate that our CstNet achieves state-of-the-art performance on both accuracy and robustness. Comprehensive ablation studies further validate the effectiveness of constraints, SurfaceKNN, and network backbone.
@@
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from a functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing procedure, to achieve a deeper understanding of CAD shapes.
Supplementary Material
F Overview
This supplementary material provides additional details that supplement the main paper and includes more experimental results.
In Sec. G, we present the details of network architecture and data preparation in validation experiments. In Sec. H, we provide the data construction and static of Param20K dataset. Finally, Sec. I offers data process for visualization and more results of constraint acquisition and classification experiments. In Sec. J, we show more details of our constraint representation and constraint learning methods.
G Validation Experiment
G.1 Constraint Expression
This section introduces the constraint expression in the validation experiments.
The constraint expression in validation experiments is defined in a simple manner: for a point, its constraint is defined by the relationship between the plane it attached and the reference plane, as shown in Fig. 11. This calculation process is carried out by openCASCADE Technology (OCCT). The constraint is represented in a one-hot encoding, where $(\mathit{1,0,0})$ indicates perpendicular, $(\mathit{0,1,0})$ indicates parallel, and $(\mathit{0,0,1})$ indicates other constraints. Each point in the point cloud is expressed as $(Coordinate,constraint)$. For example, if the coordinate of a point is $(0.1,0.2,0.3)$, and the plane it attached is perpendicular to the reference plane, it is expressed as $(0.1,0.2,0.3,\mathit{1,0,0})$, where $(0.1,0.2,0.3)$ represents the point’s coordinate, and $(\mathit{1,0,0})$ denotes the type of constraint.
Figure 11: Constraint expression in validation experiments. The red point’s constraint is expressed as the constraint type between the magenta plane and the reference plane, using one-hot encoding.
G.2 Dataset Generation
This section introduces the dataset generation procedures in the validation experiments.
The dataset generation for the validation experiments consists of the following steps:
  1. 1.
    Building the parametric templates for cuboids and prisms (with angles of 50°50°50  °, 60°60°60  °, 70°70°70  °, 80°80°80  °, 82°82°82  °, 85°85°85  °, 87°87°87  °, 89°89°89  °). Afterwards determining the dimension range for each parameter. The parametric templates and the dimensions are shown in Fig. 12.
  2. 2.
    Assigning random values to the parameters in parametric templates. Each parametric template is instantiated into 5,000 B-Rep format CAD shapes, which are stored in STEP files [35].
  3. 3.
    Using OCCT to read the STEP files and convert them into meshes. Poisson disk sampling is then applied to sample points from meshes.
  4. 4.
    For the sampled points, OCCT is used to calculate the plane to which each point is attached. The constraint type for each point is then determined based on the relationship between the attached plane and the reference plane.
  5. 5.
    Rotating the generated point clouds randomly along the XYZ axes, within a range of [25°-25°- 25  °, 25°25°25  °].
  6. 6.
    Assigning the generated point clouds to each experiments. The validation experiments involve binary classification of cuboids and prisms with specific angles. The point clouds used in each of the eight experiments are as follows (80 % for training, and 20 % for testing):
    • Experiment 1: Cuboid $\times$ 5000, Prism 50°50°50  ° $\times$ 5000
    • Experiment 2: Cuboid $\times$ 5000, Prism 60°60°60  ° $\times$ 5000
    • Experiment 3: Cuboid $\times$ 5000, Prism 70°70°70  ° $\times$ 5000
    • Experiment 4: Cuboid $\times$ 5000, Prism 80°80°80  ° $\times$ 5000
    • Experiment 5: Cuboid $\times$ 5000, Prism 82°82°82  ° $\times$ 5000
    • Experiment 6: Cuboid $\times$ 5000, Prism 85°85°85  ° $\times$ 5000
    • Experiment 7: Cuboid $\times$ 5000, Prism 87°87°87  ° $\times$ 5000
    • Experiment 8: Cuboid $\times$ 5000, Prism 89°89°89  ° $\times$ 5000
Figure 12: Parametric templates for prisms and cuboids. $a,b,c\in[0.3,2.3]$, for prisms: $b>\frac{c}{tan\alpha}$.
G.3 Model Design
This section introduces the structure of the constraint-aware model used in the validation experiments.
The constraint-aware model first employs a PointNet++ [37] backbone to predict point-wise constraints. These constraints are then concatenated with the point coordinates and passed through another PointNet++ backbone to determine the point cloud category. The model’s loss function is the sum of the classification loss and the constraint loss, as shown in Fig. 13
Figure 13: Constraint-aware model structure.
H Param20K Dataset
H.1 Why build up the Param20K dataset?
The limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes. Most existing CAD shape datasets consist of mesh files, such as MCB [18] and ESB [14]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data [21, 61] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [20] contains a large number of B-Rep files, it remains unlabeled.
H.2 Dataset Construction
This section provide details of purification process and data acquisition of Param20K dataset.
Purification: The purification process involved removing broken, duplicated, or overly complex CAD shapes requiring more than 4,000 points for point cloud representation.
Point cloud: The point cloud data in Param20K dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
H.3 Dataset Statics
This section presents the shapes count and proportion statistics for each category in the Param20K dataset. Additionally, shapes from each category are showcased.
Param20K dataset is divided into 80 % for training and 20 % for testing.
The class distribution is shown in Fig. 14.
The proportion of shapes in each category are shown in Fig. 15.
Shapes instantiated from the parametric templates designed by our team names begin with ”Num”.
Example shapes from each category are shown in Fig. 16.
Figure 14: Quantity distribution of Param20K dataset.
Figure 15: Shapes proportion of Param20K dataset.
Figure 16: Example shapes for each category in Param20K dataset.
I Constraint Acquisition and Classification Experiments
I.1 Process to Visualize Main Axis Direction
The main axis direction of a point, denoted as $v=(x,y,z)$, is visualized as $R=\frac{x+1}{2}$, $G=\frac{y+1}{2}$, and $B=\frac{z+1}{2}$. Since the main axis direction is represented as a unit vector, $-1\leqslant x,y,z\leqslant 1$, which results in $0\leqslant\frac{x+1}{2},\frac{y+1}{2},\frac{z+1}{2}\leqslant 1$.
I.2 Constraint Visualization of CstBRep
More constraints extracted by the CstBRep module from the B-Rep are visualized in Fig. 17.
Figure 17: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PMT: Primitive Type, MAD: Main Axis Direction.
I.3 Constraint acquisition on MCB
We evaluated the pre-trained CstPnt module and ParSeNet [41], HPNet [58] on the CAD shape dataset MCB [18] to predict constraints. For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison, therefore no ground truth presented in Fig. 18, and no numerical results are available. Based on the results visualized in Fig. 18, CstPnt demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
Figure 18: Constraint prediction on MCB.
I.4 Classification on MCB
We conducted evaluations on the CAD shape dataset MCB [18], with results presented in Tab. 5, from which our CstNet achieved the highest scores across all metrics, conclusions similar with experiments on Param20K dataset.
Table 5: Classification results on MCB. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
Method Acc Acc* F1 mAP
PointCNN [27] 93.89 81.85 83.86 90.13
PointNet [36] 86.78 67.70 86.55 74.08
PointNet++ [37] 87.45 73.68 88.32 91.33
SpiderCNN [57] 93.59 79.70 81.30 86.64
PointConv [51] 93.25 80.24 71.31 82.19
DGCNN [46] 92.54 74.47 76.12 74.27
3DGCN [29] 93.71 78.71 84.59 84.35
Ours 96.87 89.21 89.85 93.17
J Constraint Learning Method Details
J.1 Defects of General Constraint Representation
The simple constraint expression in validation experiments is insufficient for a comprehensive definition, as selecting an appropriate reference plane for complex CAD shapes is challenging. Moreover, traditional constraint representations suffer from non-uniqueness, as shown in Fig. 19, defining the relations between three planes involves multiple constraint options, therefore, this approach is not well-suited for deep learning models.
Figure 19: Non-uniqueness of traditional constraint expression. Left: Double vertical constraints, right: Vertical and parallel constraints.
The constraint is defined as a graph representation in some circumstances, where each node represents a primitive, and the edge between nodes represent their constraints. However, this approach is not well-suited for point cloud analysis. First, in cases where only point clouds are provided as input, the number of primitives is hard to obtain. Second, due to the unordered nature of point cloud, it is difficult to associate points with the corresponding primitives, making label design very challenging. Additionally, the graph representation essentially records the constraint relationships between each pair of primitives, for the CAD shapes contain many primitives, recording the relative positions between every pair of primitives would result in an overwhelming amount of data, making the expression of these relative positions impractical.
J.2 Unique Process of MAD
This section provides the process of Main Axis Direction and example constraint representation.
The main axis direction is processed to ensure its uniqueness, as one axis can determine two unit vectors with opposite direction. For a given axis, we first randomly select one of the two vectors, denoted as $v(x,y,z)$, and then process it using the following algorithm:
Data: Unit vector $v(x,y,z)$
Result: Unit vector after direction unification
1 if $z<0$ then
2       $v=-1\times v$;
3       Return $v$;
4
5else if $z==0$ then
6       if $y<0$ then
7             $v=-1\times v$;
8             Return $v$;
9
10      else if $y==0$ then
11             if $x<0$ then
12                   $v=-1\times v$;
13                   Return $v$;
14
15
16
return $v$
A point with constraint is expressed as Fig. 20.
Figure 20: Expression of a point with constraint.
J.3 CstBRep Module
This section explains why it is necessary to compute valid edges in the CstBRep module.
The CstBRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
Step 1: Point cloud generation (optional).
Step 2: Calculation of valid edges and the distance from each point to these valid edges. If the minimum distance is below a threshold, the point is classified as a edge-near point.
Step 3: Calculation of the primitives to which each point attached, followed by analysis to determine point’s Primitive Type and the Main Axis Direction.
In Step 2, valid edges are computed, as this helps to reduce the potential for data to mislead the deep learning model. The definition of valid edge is illustrated in Fig. 21. This definition is based on the observation that a complete cylinder in the B-Rep data stored in STEP files is often split into two half-cylinder faces. As a result, two generatrices form edges on the cylindrical surface, but the deep learning model cannot predict these edges, as shown in Fig. 22. Furthermore, in other cases, edges corresponding to smoothly connected faces are also difficult to identify, leading to confusion for deep learning models.
Figure 21: Zebra stripes of primitives corresponding to invalid and valid edges. Left: G1 or higher continuous at the invalid edges, right: G0 continuous at the valid edges.
Figure 22: Cylinder in B-Rep representation. A complete cylindrical surface is represented as a combination of two half-cylinder surface, resulting in edges along the generatrices that cannot be recognized, thereby causing confusion for deep learning model.
J.4 CstPnt Module
This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CstPnt Module.
The CstPnt Module is used to predict constraints from point clouds. The design of the CstPnt Module is inspired by the following calculation process:
For a point $\boldsymbol{p}$ in point cloud, its constraint could be calculated by the following steps:
  1. 1.
    Identify neighbor points around $\boldsymbol{p}$.
  2. 2.
    From all neighbor points, identify those that belong to the same primitive as $\boldsymbol{p}$; these points are referred to as valid points.
  3. 3.
    Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
  4. 4.
    Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
In these steps, it is necessary to compute the valid points. We define the valid point as the neighbors that lie on the same primitive as the center point $\boldsymbol{p}$, as visualized in Fig. 23. These valid points can be used to determine whether the center point is near an edge, as well as to calculate the main axis direction and primitive type of the attached primitive, while the invalid points do not contribute to these calculations.
Figure 23: Valid and invalid points for constraint calculation.
J.5 SurfaceKNN
This section introduces the algorithm of SurfaceKNN.
SurfaceKNN based on the assumption that when KNN is applied with a small number of neighbors, the output points are searched along the shape’s surface. This assumption has been validated in most cases. Leveraging this assumption, SurfaceKNN is accomplished by applying KNN with small number of neighbors iteratively. During each iteration, KNN is used to search a small set of neighboring points, which then serve as new central points for subsequent searches. This process continues until the desired number of points is obtained, as illustrated in Fig. 24.
Figure 24: Algorithm of SurfaceKNN.

### v6

@@ 3.4 Constraint-Aware Feature Learning
The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD-Adj-PT separately, and attention between features is then applied to update constraint feature, with details in Fig. 7.
@@ 4.1 Experimental Settings
Param20K dataset: Param20K contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes BRep data, mesh, and point cloud. All BRep data are collected from the following three sources. More details in the supplementary. Sources 1 (36.05%): Downloaded and purified from TraceParts [20]. The CAD shapes on TraceParts are collected from manufacturing companies. Sources 2 (24.12%): Provided by a CAD company [1] that collaborates with our team. These shapes are all mechanical components. Sources 3 (39.83%): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
@@ 4.2 Constraint Acquisition
Extracting MAD-Adj-PT from BRep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig. 8. By referring to the model row, CstBRep performs satisfactorily in both MAD-Adj-PT extraction. Predicting MAD-Adj-PT from point cloud is achieved by the CstPnt module. The training set is sourced from 25 trunks of ABC [22], each trunk containing 10,000 BRep files. We utilized the CstBRep to extract constraints as Ground Truth for model training. After the training finished, the MSE loss of MAD is 0.0559, the prediction accuracy of Adj and PT reached 94.62% and 94.23%, respectively. In subsequent experiments, the above pre-trained CstPnt is utilized to predict MAD-Adj-PT with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods when comparing, even when BRep data is available, the CstBRep module will not be activated to obtain MAD-Adj-PT.
After CstPnt was pre-trained, we evaluated it on a ABC new trunk and Param20K dataset. The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available datasets in $Suppl.$ Since there is no existing model that directly predicts MAD-Adj-PT, we modified the published CAD-specific methods [63, 43] by adding MLPs as prediction head to enable these predictions. The experiment results are shown in Tab. 1, from which our CstPnt achieves the highest accuracy. Furthermore, from ABC to Param20K, all methods’ performance decreased, but our method dropped least accuracy. ParSeNet and HPNet may be negatively affected by global features, which impaired their performance on unseen datasets. Fig. 8 presents the visualization of constraint prediction results, which shown that our method is closest to the ground truth. Furthermore, from ABC to Paeam20K, our method demonstrates excellent generalization performance on unseen datasets. In contrast, while ParSeNet and HPNet exhibit acceptable prediction results on the ABC new trunk, they performed unsatisfactory on the unseen Param20K dataset.
@@
Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in $Suppl.$
@@ 4.3 Classification
To validate the ability of our CstNet in identifying CAD shapes that are visually similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in Sec. 3.1. Results shown in Fig. 9, in which our CstNet achieved the best instance accuracy in distinguishing 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrates the strong ability of CstNet in discriminating CAD shapes with similar appearance but constraints different. For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, which is attributed to the differences between them being so subtle that no one is capable of distinguishing them.
@@
The classification of prisms and cuboids is predominantly influenced by constraints, while the comprehension of common CAD shapes necessitates the integration of both constraint and geometric features. CstNet achieves this by concatenating xyz with MAD-Adj-PT in Stage 2, thereby integrating geometric features into constraint features. To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset, more results on other public available datasets in $Suppl.$ Results are presented in Tab. 2, where our CstNet outperformed other methods across both metrics, validated the effectiveness of utilizing constraint and geometric features.
@@
To access the robustness of CstNet, we evaluated it under rotated point clouds. Since point clouds are pre-processed by aligning their centroids to the origin and scaling their ranges into a unit cube, the performance of all methods is unaffected by point cloud shifts and scaling. Comparison results are depicted in Fig. 10, from which our CstNet demonstrates strongest robustness to point cloud rotation. The dashed line in Fig. 10 show that the performance of CstPnt module remains nearly unaffected by rotation, owing to it learned various local features from ABC dataset.
@@ 4.4 Ablation Studies
Our CstNet leverages all three components of constraint representation: MAD-Adj-PT. Additionally, we utilized the SurfaceKNN to obtain more efficient neighbors for each point. To validate the effectiveness of our design, we conducted a series of ablation studies on Param20K dataset.
For each component of constraint representation, we evaluated their importance by disabling one or all at a time. This operation is performed by setting the disabled components as a zero tensor. Classification results are presented in Tab. 3, which show that all three components contribute to improving CstNet’s performance. To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 3, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
@@
To validate the effectiveness of Stage 2 backbone, we conducted a comprehensive evaluation by feeding (x, y, z, MAD, Adj, PT) into multiple classification backbone for comparison. The results are presented in Tab. 4, from which CstNet demonstrates best performance under both predicted and label MAD-Adj-PT, indicating that our design achieves the best utilization of MAD-Adj-PT. Furthermore, although the accuracy of using predicted MAD-Adj-PT is lower than that of using label MAD-Adj-PT, it still shows a significant improvement compared to the baselines without MAD-Adj-PT. This highlights the effectiveness of the predicted MAD-Adj-PT in improving model performance.
Table 4: Effectiveness of Stage2 Backbone. Upper half: predicted constraint representation from CstPnt, lower: label constraint representation.
@@ 5 Conclusions
In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly representation. Afterward, the CstNet is proposed for extracting and leveraging constraint representation. Finally, we built a multi-modal CAD dataset Param20K. Extensive experiments demonstrate that the CstNet achieves SOTA performance on both accuracy and robustness. Comprehensive ablation studies further validate the effectiveness of the proposed constraint representation, SurfaceKNN, and network backbone.
@@
CAD shapes possess numerous attributes. We analyze the constraints from a functional perspective in this work. In future work, we plan to investigate additional aspects such as part design and manufacturing procedure, to achieve a deeper understanding of CAD shapes.
Acknowledgements
This work was supported by National Key Research and Development Program of China (Grant No.2022YFB3303101).